An 8-digit no. is a multiple of 73 and 137. If the second digit from left is 7 then 6th digit from left of the n. is ?

Grade:11

3 Answers

Sher Mohammad IIT Delhi
10 years ago
73*137=10001, whenever it will be multiplied by any digit number , will give 8 digit number, let 3 digit number is pqrs
10001*pqr=pqrs*0000+pqrs=pqrs0pqrs, see second and 6 digit is same, so 6th digit will be 7.

sher mohammad
b.tech, iit delhi
lokesh niranjan
9 years ago
Hi sher mohamed,
Please explain the answer in brief.
lokesh
Vaishali Gaur
6 years ago
73 and 137 are prime number.73×137=10001When we multiply 10001 by (1700 to 1799)We get 8 digit number as 10001×1700=1700170010001×1701=17011701 etc.These all digit numbers are multiples of 73 and 137 Second digit from right is 7Hence,sixth from the left is 7.